OpenStudy (anonymous):

what is the x-intercept of the line represented by y=2x+6? A)-3 B)2 c)3 D)6

OpenStudy (campbell_st):

the slope intercept form of a line is y = mx + b m = slope, b = y-intercept hope this helps you find the intercept for the line y = 2x + 3

OpenStudy (anonymous):

make the y=0 again and solve for x

OpenStudy (campbell_st):

oops its y = 2x + 6

OpenStudy (anonymous):

mistake lol like me

OpenStudy (anonymous):

can you please help me on how to solve it step by step if guys know

OpenStudy (anonymous):

PLEASE >_<

OpenStudy (thomas5267):

I have just did a similar question here. http://openstudy.com/study#/updates/4f866164e4b0505bf0864773

OpenStudy (anonymous):

0=2x+6 -6=2x -3=x this should be the x intercept if i'm not wrong

OpenStudy (anonymous):

ok thank you both

OpenStudy (anonymous):

no problem any time(when i'm on)